Solve the trigonometric equation:3cos2theta - sin theta = 2

Grade:10

1 Answers

sonika p
6 years ago
3cos2theta – sin theta = 2
3 (1 – 2sin^2 theta) -sin theta =2
3 – 6sin^2 theta – sin theta =2
1 – 6sin^2 theta – sin theta =0
6sin^2 theta + sin theta – 1 =0
sin theta = ( -1 +/- root(1+24))/12
sin theta = -6/12, sin theta = 4/12
sin theta = -1/2 or sin theta = 1/3
theta = -pi/6 or 19.47o